### v1

@@ 4.1 Experimental Details
1) Datasets: Five simulated datasets, including Pavia University cite ?, [7, Pavia Center cite ?, [7, Indian Pines cite ?, [4, Washington DC cite ?, [48, 47 and University of Houston cite ?, [20 were used for simulations in our experiments. The five synthetic datasets, each with a simulated PSF and SRF, are used to generate two observed images: LR-HSI and HR-MSI. More specifically, we utilize a Gaussian blur of $5\times 5$ pixels with a 0 mean and 1 standard deviation to simulate the PSF for generating the LR-HSI. The downsampling ratio was set to 4 for all datasets. The spectral response of the IKONOS satellite (a Nikon D700 camera) was used to simulate the SRF for generating the HR-MSI. The i.i.d Gaussian noise was added to HR-MSI and LR-HSI with signal-to-noise ratios (SNRs) of 30dB, respectively. The details for the five datasets are summarized as follows.
@@
2) Comparison methods: Since our method is an unsupervised algorithm, we mainly compare it to unsupervised fusion methods; such a choice also aligns with the practical demands of real-world scenarios. Specifically, we compare the proposed method with several SOTA unsupervised fusion methods, including CNMF cite ?, [43, HySure cite ?, [28, HyCoNet cite ?, [46, CNN-FUS cite ?, [10, MSE-SURE cite ?, [24, and E2E-FUS cite ?, [39. For baselines: CNMF, HySure, CNN-FUS, and E2E-FUS methods are implemented using MATLAB R2023a on Windows 11 with 32GB RAM, while HyCoNet, MSE-SURE, and CLoRF are evaluated on an RTX 4090 GPU with 32GB RAM. Besides, parameters in baselines are manually fine-tuned to achieve optimal results in all experiments.
@@
4) Hyperparameters of CLoRF: For all tasks, both the spatial-INR and spectral-INR adopt the SIREN network with an initial network parameter $\omega_{0}=30$. The spatial-INR has 5 hidden layers with a size of 512, while the spectral-INR has 2 hidden layers with a size of 128 (In Indian Pines, the spatial-MLP has 3 hidden layers with a size of 512, while the spectral-MLP also has 3 hidden layers with a size of 256). The learning rate is set to 3e-5, and the training epochs are fixed to 30000. An early stopping strategy is implemented to prevent overfitting for all datasets. Furthermore, the hyperparameters are summarized in Table 1.
@@
Hyperparameter Pavia University Pavia Center Indian Pines Washington DC University of Houston
$K$ 9 9 22 11 10
$\lambda$ 1.25 1.25 0.55 1.25 1
$\eta$ 0.0025 0.0050 0.0060 0.0025 0.0025
Table 2: Quantitative performance comparison with different algorithms on the different datasets. The best results are bold-faced, and runner-ups are underlined. (MPSNR ↑, MSSIM ↑, SAM ↓, ERGAS ↓)
@@
4.2 Experimental Results
@@
The fusion performance of all methods on five datasets, evaluated across MPSNR, MSSIM, SAM, and ERGAS metrics, is shown in Table 2. CLoRF consistently outperforms others in terms of quantitative results across most scenarios, affirming the efficacy of the HSI-MSI fusion task. Moreover, CLoRF exhibits commendable performance in spatial structures. our method outperforms E2E-FUS by 2.19% in terms of MPSNR. Nonetheless, in certain datasets, particularly those with a high number of bands, CLoRF slightly lags behind MSE-SURE in SAM index performance. It is notable that MSE-SURE, throughout its training phase, is equipped with knowledge of noise levels in LR-HSI and HR-MSI, and leverages back-projection techniques to enhance spectral detail capture, while our method does not require this information.
@@
To assess the effectiveness of various methods in preserving spatial structures, Fig. 3-7 illustrates estimated HR-HSI on five simulated datasets. The first rows exhibit fusion results virtually, showcasing the 1st band of Pavia University, the 8th band of Pavia Center, the 160th band of Indian Pines, the 5th band of Washington DC Mall, and the 46th band of the University of Houston. Conversely, the second rows depict corresponding error maps, illustrating the mean squared error (MSE) between GT and the estimated HR-HSI. While the results for the images in the first rows appear similar across most methods, with differences almost imperceptible to the naked eye, notable disparities emerge in the error maps depicted in the second column. It is evident that CLoRF exhibits minimal noise and preserves superior spatial structures. As shown in Fig. 8, we plot the two randomly selected spectral vectors reconstructed by CLoRF and E2E-FUS with four datasets. It is apparent that CLoRF effectively preserves the high-frequency information compared to E2E-FUS in some pixels. However, in some specific pixels, the performance of the E2E-FUS method is superior to CLoRF.
@@ 4.3 Arbitrary Resolution
Here, we evaluate the performance of CLoRF to fuse HSIs at arbitrary spatial and spectral resolutions. As shown in Fig. 1, our model is well-trained based in an unsupervised manner, and we can input any scale spatial and spectral position coordinates during the inference stage, obtaining arbitrary resolutions in both spatial and spectral domains. We use PSF and downsampling ratio of the same size as in Sec. 4.1, but different sizes of SRF (sampled from SRF in Sec. 4.1) to synthesize the reduced-resolution LR-HSI ($42\times 42\times 50$) and HR-MSI ($168\times 168\times 4$) with Pavia University for training spatial-INR and spectral-INR. Then, we predict HR-HSIs of arbitrary resolutions using different scales of spatial and spectral coordinates. For a fair comparison, our method is directly compared with bicubic interpolation from the original LR. As shown in Table 3, we can fix spectral resolution to obtain arbitrary spatial resolution, fix spatial resolution to obtain arbitrary spectral resolution, or simultaneously achieve resolutions in both spatial and spectral domains. CLoRF performs nearly as well as bicubic interpolation in spatial resolution when spectral resolution is fixed, yet it notably surpasses bicubic when spatial resolution is fixed while achieving spectral resolution. Moreover, when simultaneously enhancing resolutions in both spatial and spectral domains, CLoRF consistently outperforms bicubic interpolation. From Fig. 9-10, it can be seen that when simultaneously upsampling in both spatial and spectral domains, CLoRF obtains finer details compared to the bicubic interpolation.
@@
Table 3: Experimental results in spatial (fixed spectral), spectral (fixed spatial), and (spatial, spectral) with arbitrary resolution, metric: MPSNR/MSSIM. The original dimension of HR-HSI is (168, 168, 50). The best are highlighted.
@@ 4.4 Ablation Study
Table 4: Ablation study of the TV loss.
@@
Table 5: Ablation study of the different activation functions. (MPSNR / MSSIM)
@@
2) Different activation functions: Table 5 displays the results obtained using different activation functions, such as ReLU, ReLU+Position Encoding (PE), Gauss cite ?, [26, Spder cite ?, [27, and Sine cite ?, [29. Notably, the Sine activation function demonstrates outstanding performance across all metrics. Therefore, we have made it the default activation function for our model.
@@ 4.5 PAN-HSI Fusion
1) Dataset: The dataset acquired for the PRISMA contest1 cite ?, [33, namely RR1. The image is of size $900\times 900\times 59$. The synthetic datasets, plus a simulated PSF and SRF, were used to generate two observed images: LR-HSI and PAN. PSF is set to be the same as the 4.1. The wavelength cite ?, [33 was used to simulate the SRF to generate the PAN image. The i.i.d Gaussian noise is added to the LR-HSI (30dB) and PAN (30dB) image.
2) Compared Methods: We primarily compare CLoRF with two categories of baselines: model-based methods include GS cite ?, [19, GSA cite ?, [2, AWLP cite ?, [32, MTF-GLP cite ?, [1, and HySure cite ?, [28, while the unsupervised deep learning method is R-PNN cite ?, [14. All training parameters for CLoRF remain the same for Pavia University.
@@
Table 6: Quantitative performance comparison with different algorithms on the RR1 dataset. The best results are bold-faced, and runner-ups are underlined. (MPSNR ↑, MSSIM ↑, SAM ↓, ERGAS ↓)

### v2

@@ 4.1 Experimental Details
1) Datasets: We evaluate the performance of fusion using both synthetic and real datasets. Seven simulated datasets, including Pavia University cite ?, [9, Pavia Center cite ?, [9, Indian Pines cite ?, [4, Washington DC cite ?, [61, 60, University of Houston cite ?, [25, Peppers cite ?, [52, and Superballs cite ?, [52 were used for simulations in our experiments. The seven synthetic datasets, each with a simulated PSF and SRF, are used to generate two observed images: LR-HSI and HR-MSI. More specifically, we utilize a Gaussian blur of $5\times 5$ pixels with a 0 mean and 1 standard deviation to simulate the PSF for generating the LR-HSI. The downsampling ratio was set to 4 for all datasets. The spectral response of the IKONOS satellite (a Nikon D700 camera) was used to simulate the SRF for generating the HR-MSI. The i.i.d Gaussian noise was added to HR-MSI and LR-HSI with signal-to-noise ratios (SNRs) of 30 dB, respectively. The details for the eight datasets are summarized as follows.
@@
  • Peppers and Superballs: We utilize the widely-used CAVE dataset cite ?, [52, which is a ground-based HSI dataset commonly employed in HSI-MSI fusion. The dataset comprises 32 high-resolution HSIs, each with a size of $512\times 512\times 31$.
  • Real Data: The LR-HSI is collected by the Hyperion sensor cite ?, [50, which is of the size $120\times 120\times 89$. The HR-MSI with 13 bands is taken by the Sentinel-2A satellite. Four bands are employed for the test, and the spatial downsampling factor is 3, i.e., the size of HR-MSI is $360\times 360\times 4$. These four bands are extracted from bands 2, 3, 4, and 8, with the central wavelengths being 490, 560, 665, and 842 nm, respectively.
2) Comparison methods: Since our method is an unsupervised algorithm, we mainly compare it to unsupervised fusion methods; such a choice also aligns with the practical demands of real-world scenarios. Specifically, we compare the proposed method with several SOTA unsupervised fusion methods, including CNMF cite ?, [54, HySure cite ?, [36, HyCoNet cite ?, [59, CNN-FUS cite ?, [13, MSE-SURE cite ?, [31, and E2E-FUS cite ?, [47. For baselines: CNMF, HySure, CNN-FUS, and E2E-FUS methods are implemented using MATLAB R2023a on Windows 11 with 32GB RAM, while HyCoNet, MSE-SURE, CLoRF, and supervised methods are evaluated on an RTX 4090 GPU with 32GB RAM. Besides, parameters in baselines are manually fine-tuned to achieve optimal results in all experiments.
@@
4) Hyperparameters of CLoRF: For all tasks, both the spatial-INR and spectral-INR adopt the SIREN network with an initial network parameter $\omega_{0}=30$. The spatial-INR has 5 hidden layers with a size of 512, while the spectral-INR has 2 hidden layers with a size of 128 (In Indian Pines, the spatial-MLP has 3 hidden layers with a size of 512, while the spectral-MLP also has 3 hidden layers with a size of 256). Spatial-INR and spectral-INR are jointly optimized based on the loss function 7, using the Adam optimizer for optimization. The learning rate is set to 3e-5, and the training epochs are fixed to 30000. An early stopping strategy is implemented to prevent overfitting for all datasets. Furthermore, the hyperparameters are summarized in Table 1.
@@
Hyperparameter Pavia University Pavia Center Indian Pines Washington DC Houston Peppers Superballs
$K$ 9 9 22 11 10 10 10
$\lambda$ 1.25 1.25 0.55 1.25 1 1 1
$\eta$ 0.0025 0.0050 0.0060 0.0025 0.0025 0.0025 0.0025
Table 2: Quantitative performance comparison with different algorithms on the different datasets. The best results are bold-faced, and runner-ups are underlined. (MPSNR ↑, MSSIM ↑, SAM ↓, ERGAS ↓).
@@
Peppers MPSNR 40.00 36.21 41.15 44.86 43.25 43.23 47.65
MSSIM 0.96 0.94 0.96 0.99 0.99 0.98 0.99
SAM 11.62 10.59 4.77 7.36 6.34 7.02 4.01
ERGAS 4.92 9.04 6.29 2.65 2.87 2.94 1.94
Superballs MPSNR 43.88 39.14 43.05 45.26 44.74 45.80 46.14
MSSIM 0.97 0.96 0.98 0.98 0.98 0.98 0.98
SAM 10.84 10.40 7.98 7.47 6.84 7.42 6.40
ERGAS 3.58 5.82 5.98 2.92 3.10 2.72 2.71
@@
Figure 8: The first row shows the Peppers image (22th band) of the estimated HR-HSI, and the second row shows the error map between the estimated image and GT.
Figure 9: The first row shows the Superballs image (30th band) of the estimated HR-HSI, and the second row shows the error map between the estimated image and GT.
@@
4.2 Evaluation on Synthetic Data
Here, we consider synthetic data and evaluate proposed CLoRF under various scenarios.
@@ 4.2 Evaluation on Synthetic Data
1) HSI-MSI fusion: The fusion performance of all methods on seven datasets, evaluated across MPSNR, MSSIM, SAM, and ERGAS metrics, is shown in Table 2. CLoRF consistently outperforms others in terms of quantitative results across most scenarios, affirming the efficacy of the HSI-MSI fusion task. Moreover, CLoRF exhibits commendable performance in spatial structures. our method outperforms E2E-FUS by 2.19% in terms of MPSNR. Nonetheless, in certain datasets, particularly those with a high number of bands, CLoRF slightly lags behind MSE-SURE in SAM index performance. It is notable that MSE-SURE, throughout its training phase, is equipped with knowledge of noise levels in LR-HSI and HR-MSI, and leverages back-projection techniques to enhance spectral detail capture, while our method does not require this information.
@@
To assess the effectiveness of various methods in preserving spatial structures, Fig. 3-9 illustrates estimated HR-HSI on seven simulated datasets. The first rows exhibit fusion results virtually, showcasing the 1st band of Pavia University, the 8th band of Pavia Center, the 160th band of Indian Pines, the 5th band of Washington DC Mall, the 46th band of the University of Houston, the 22nd band of the Peppers, and the 30th band of the Superballs. Conversely, the second rows depict corresponding error maps, illustrating the mean squared error (MSE) between GT and the estimated HR-HSI. While the results for the images in the first rows appear similar across most methods, with differences almost imperceptible to the naked eye, notable disparities emerge in the error maps depicted in the second column. It is evident that CLoRF exhibits minimal noise and preserves superior spatial structures. As shown in Fig. 10, we plot the two randomly selected spectral vectors reconstructed by CLoRF and E2E-FUS with four datasets. It is apparent that CLoRF effectively preserves the high-frequency information compared to E2E-FUS in some pixels. However, in some specific pixels, the performance of the E2E-FUS method is superior to CLoRF.
2) Runtime and complexity analysis: We have conducted a runtime and complexity analysis of various methods on the Houston dataset, with the experimental results reported in Table 3. Please note that the runtime for supervised learning methods includes both training and inference time. Plug-and-play methods require pre-trained networks. In comparison to self-supervised learning methods, CLoRF demonstrates a significant advantage in terms of runtime efficiency and GFLOPs. Additionally, CLoRF excels in running time and model parameters when compared to supervised learning baselines.
Table 3: Runtime, parameters, and GFLOPs of all methods for Houston dataset. The best results are bold-faced.
Methods Times (s) Paras (M) GFLOPs
Model-based CNMF 5.70 (CPU) / /
HySure 57.60 (CPU) / /
Plug-and-Play CNN-FUS 114.06 (CPU) / /
E2E-FUS 107.00 (CPU) / /
Deep-learning HyCoNet 1332.38 (GPU) 0.58 152.63
MSE-SURE 478.23 (GPU) 1.10 149.92
(Self-supervised) CLoRF 446.86 (GPU) 1.36 121.75
Deep-learning ConSSFCNN 473.51 (GPU) 4.79 78.37
ResTFNet 525.07 (GPU) 23.24 84.73
(Supervised) SSR-Net 504.51 (GPU) 2.87 46.69
MCT-Net 1128.71 (GPU) 36.09 117.86
3) Performance on different downsampling ratios: We evaluate the model’s performance when there is a significant difference in resolution between the input LR-HSI and the desired HR-HSI. Specifically, we choose the downsampling ratio in $\{4,8,16\}$. The experimental results are shown in Table 4, where we observe that our method still demonstrates its advantage when there is a large resolution gap between the LR-HSI and HR-HSI.
Table 4: Performance on different downsampling ratios with 30 dB noise for Pavia University. The best results are bold-faced.
Ratios Methods MPSNR MSSIM SAM ERGAS
4 CNMF 32.70 0.92 3.55 3.63
HySure 32.86 0.93 5.81 3.71
HyCoNet 40.10 0.97 3.10 1.41
CNN-FUS 41.57 0.98 2.38 1.38
MSE-SURE 41.31 0.98 2.20 1.44
E2E-FUS 40.92 0.98 2.31 1.50
CLoRF 42.23 0.99 2.05 1.31
8 CNMF 24.23 0.74 7.67 4.82
HySure 23.87 0.74 12.55 4.88
HyCoNet 39.67 0.97 3.18 0.73
CNN-FUS 40.09 0.97 2.76 0.85
MSE-SURE 40.71 0.98 2.40 0.87
E2E-FUS 40.29 0.98 2.61 0.83
CLoRF 41.56 0.98 2.29 0.69
16 CNMF 21.27 0.62 12.32 3.50
HySure 17.54 0.58 21.06 4.67
HyCoNet 39.47 0.97 3.25 0.37
CNN-FUS 40.08 0.98 2.70 0.42
MSE-SURE 39.63 0.98 2.92 0.44
E2E-FUS 40.21 0.98 2.63 0.42
CLoRF 41.11 0.98 2.33 0.37
4) Comparison with supervised methods: To provide a broader context for the performance of the proposed approach, we compare it with recent four supervised methods, such as ConSSFCNN cite ?, [19, ResTFNet cite ?, [27, SSR-Net cite ?, [58, MCT-Net cite ?, [46. For fairness, since our approach is self-supervised, we evaluate it on the same test sets of the datasets used by the supervised methods, as shown in Table 5. Our method demonstrates comparable performance to MCT-Net and even outperforms the supervised methods on the Pavia Center dataset.
Table 5: Performance of different supervised methods on two datasets, and the downsampling ratio is 4 with 30 dB noise. The best results are bold-faced.
5) Performance under degeneration with estimated PSF and SRF: To test the impact of degradation operators in the spatial and spectral domains on HR-HSI fusion results, we assume that the degradation operators are unknown in the simulated dataset, and we estimate these operators using the method suggested in cite ?, [36. In Table 6, we present the results of both semi-blind and fully-blind experiments. Note that the CNMF and HySure methods are fully blind in the experiment. They are not included in the semi-blind experiments for comparison. As shown in Table 6, blind fusion is more challenging than semi-blind fusion, all methods perform similarly underestimated degradations in the spatial domain. However, in blind fusion, the performance of all methods decreases to some extent. Nevertheless, our method still demonstrates greater robustness in comparison to other approaches.
Table 6: Performance under generation produced by semi-blind (estimate PSF) and blind (estimate both PSF and SRF) with 30 dB noise and the downsampling ratio is 4 for Pavia University. The best results are bold-faced.
Methods MPSNR MSSIM SAM ERGAS
Semi-blind HyCoNet 40.10 0.97 3.10 1.41
CNN-FUS 40.04 0.97 2.55 1.64
MSE-SURE 41.11 0.98 2.20 1.49
E2E-FUS 40.77 0.98 2.34 1.52
CLoRF 41.40 0.98 2.32 1.40
Blind CNMF 32.70 0.92 3.55 3.63
HySure 32.86 0.93 5.81 3.71
HyCoNet 38.06 0.95 3.97 1.83
CNN-FUS 38.32 0.96 3.02 2.12
MSE-SURE 36.14 0.96 4.36 2.90
E2E-FUS 31.33 0.84 7.53 5.34
CLoRF 38.98 0.97 3.00 1.88
@@ 4.3 Arbitrary Resolution
Here, we evaluate the performance of CLoRF to fuse HSIs at arbitrary spatial and spectral resolutions. As shown in Fig. 1, our model is well-trained based in an unsupervised manner, and we can input any scale spatial and spectral position coordinates during the inference stage, obtaining arbitrary resolutions in both spatial and spectral domains. We use PSF and downsampling ratio of the same size as in Sec. 4.1, but different sizes of SRF (sampled from SRF in Sec. 4.1) to synthesize the reduced-resolution LR-HSI ($42\times 42\times 50$) and HR-MSI ($168\times 168\times 4$) with Pavia University for training spatial-INR and spectral-INR. Then, we predict HR-HSIs of arbitrary resolutions using different scales of spatial and spectral coordinates. For a fair comparison, our method is directly compared with bicubic interpolation from the original GT. As shown in Table 7, we can fix spectral resolution to obtain arbitrary spatial resolution, fix spatial resolution to obtain arbitrary spectral resolution, or simultaneously achieve resolutions in both spatial and spectral domains. CLoRF performs nearly as well as bicubic interpolation in spatial resolution when spectral resolution is fixed, yet it notably surpasses bicubic when spatial resolution is fixed while achieving spectral resolution. Moreover, when simultaneously enhancing resolutions in both spatial and spectral domains, CLoRF consistently outperforms bicubic interpolation. From Fig. 11-12, it can be seen that when simultaneously upsampling in both spatial and spectral domains, CLoRF obtains finer details compared to the bicubic interpolation. Furthermore, we train a smaller size of data (LR-HSI ($42\times 42\times 50$) and HR-MSI ($168\times 168\times 4$) with Pavia University for training spatial-INR and spectral-INR) and infer the desired HR-HSI with spatial super-resolution factor in $\{2,4,8,16\}$ and a fixed spectral upsampling resolution 93. Figure 13 demonstrates that the trained CLoRF model successfully super-resolves the original HR-HSI $(168\times 168\times 50)$ to arbitrary resolutions with a large range of flexibility. However, at a super-resolution factor of 16, the image exhibits slight fluctuations within some regions, which shows the inherent limitations.
@@
Figure 13: The original resolution of HR-HSI is $(168,168,50)$, the spatial super-resolution factor in $\{2,4,8,16\}$. Visualize the 40th band at different resolutions for Pavia University.
Table 7: Experimental results in spatial (fixed spectral), spectral (fixed spatial), and (spatial, spectral) with arbitrary resolution, metric: MPSNR/MSSIM. The original dimension of HR-HSI is (168, 168, 50). The best results are bold-faced.
@@ 4.4 Ablation Study
Table 8: Ablation study of the TV loss. The best results are bold-faced.
@@
Table 9: Ablation study of the different activation functions. (MPSNR / MSSIM). The best results are bold-faced.
@@
2) Different activation functions: Table 9 displays the results obtained using different activation functions, such as ReLU, ReLU+Position Encoding (PE), Gauss cite ?, [34, Spder cite ?, [35, and Sine cite ?, [37. Notably, the Sine activation function demonstrates outstanding performance across all metrics. Therefore, we have made it the default activation function for our model.
@@
4.5 Evaluation on Real Dataset
To further illustrate the effectiveness of the proposed method, we conduct a real-world fusion experiment. It is important to note that the ground-truth spatial and spectral degradation operators are not available for real data. Consequently, we estimate these operators using the approach suggested in cite ?, [36. For parameter selection in real data, we first apply the classical noise estimation algorithm proposed in cite ?, [5 to obtain a preliminary estimate of the noise intensity of the LR-HSI, which is SNR=32.75 dB. Since the estimated noise intensity is close to SNR = 30 dB, In our method, we set $K$ and $\eta$ to be consistent with Pavia University, and $\lambda=1.8$. The experimental parameters for the other methods remain consistent with those of the Pavia University dataset.
Since there is no ground-truth for the HR-HSI, we visualize the estimated HR-HSIs of all methods in Fig. 15 together with the image quality score measured by a non-reference image quality metric cite ?, [51. We find that the proposed CLoRF recovers more details and obtains the best image quality.
Figure 15: Visual the real data (14th band) of the estimated HR-HSI, evaluated with a no-reference hyperspectral imaging quality score for various methods applied to the real HypSen dataset in blind fusion tasks.
@@ 4.6 PAN-HSI Fusion
1) Dataset: The dataset acquired for the PRISMA contest1 cite ?, [41, namely RR1. The image is of size $900\times 900\times 59$. The synthetic datasets, plus a simulated PSF and SRF, were used to generate two observed images: LR-HSI and PAN. PSF is set to be the same as the 4.1. The wavelength cite ?, [41 was used to simulate the SRF to generate the PAN image. The i.i.d Gaussian noise is added to the LR-HSI (30 dB) and PAN (30 dB) image.
2) Compared Methods: We primarily compare CLoRF with two categories of baselines: model-based methods include GS cite ?, [24, GSA cite ?, [2, AWLP cite ?, [40, MTF-GLP cite ?, [1, and HySure cite ?, [36, while the unsupervised deep learning method is R-PNN cite ?, [17. All training parameters for CLoRF remain the same for Pavia University.
@@
Table 10: Quantitative performance comparison with different algorithms on the RR1 dataset. The best results are bold-faced, and runner-ups are underlined. (MPSNR ↑, MSSIM ↑, SAM ↓, ERGAS ↓).